# RoToR: Towards More Reliable Responses for Order-Invariant Inputs

### v1

@@ 4.1 Baselines
Original causal LM with no modifications (Orig.) were compared, which processes text sequentially. Also, we compare RoToR against previous zero-shot order-invariant LMs discussed in Sec. 3.1, namely PCW Ratner et al. (2023), PINE Wang et al. (2024), and Set-Based Prompting McIlroy-Young et al. (2024). We use the LLaMA 3.1 AI (2024) 8B-Instruct444meta-llama/Meta-Llama-3.1-8B-Instruct as our backbone model for all of our experiments [^5] Results on 70B model is available in Appendix, Tab. 5.. As our method doesn’t need training, a single A6000 GPU was sufficient to run all of the experiments with our primary backbone model.
@@ 4.2 Benchmarks with listwise inputs
We experiment with three benchmarks involving real-world listwise input data. Examples of exact inputs and outputs are provided in Appendix A.7. All reported scores are rounded to the nearest tenth.
Knowledge Graph Question Answering (KGQA)
In KGQA tasks, the model takes facts over knowledge graphs represented as (subject, relation, object), and answers the given question based on the given facts. We basically follow the KGQA dataset preprocessing and evaluation setup from  Baek et al. (2023), which uses WebQSP Yih et al. (2016) modified from  Berant et al. (2013) and Mintaka Sen et al. (2022) with Wikidata for knowledge source, and use the Exact Match (EM), Accuracy (Acc), and F1 score metric for evaluation. We also use MPNet Song et al. (2020) as a dense retriever to retrieve top-k facts over each question, and experiment with segment size of 30 and 50. Replication details and example dataset format are at Appendix Sec. A.3, Fig. 9 and Fig. 10, respectively. Along with measuring the performance of the initial input ordering, we report performance after we shuffle the order of segments with 3 different seeds to see shuffle robustness.
Lost in the middle (LitM) We use the Lost in the Middle (LitM) benchmark Liu et al. (2024), which draws from 2655 queries in the Natural Questions (NQ) dataset. It provides sets of (10, 20, 30) passages, placing the gold passage at predetermined positions (e.g., 0, 4, 9) and filling the remaining slots with irrelevant passages, testing the model’s robustness to passage ordering. Following  Liu et al. (2024), we measure answer quality with the best_subspan_em metric. Experiments on LitM found that eliminating the effect of index bias is another important detail for invariant models to be effective: (Appendix Sec. B) thus, we also experiment with a variant of the benchmark with index bias eliminated (prompt at Appendix Fig. 8) in order to further investigate the impact with and without index bias.
MMLU The Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2021) consists of 57 diverse sub-tasks with a total of 14,015 queries to measure general performance of LMs about the knowledge of the world. Despite its popularity, many works report that performance fluctuates heavily depending on the order of choices Gupta et al. (2024); Pezeshkpour and Hruschka (2023); Wei et al. (2024); Alzahrani et al. (2024); Zheng et al. (2024a) and is widely investigated to measure the positional bias of the model. On experiments on MMLU, we notice that a lot of proportions consist of ordering-sensitive inputs, which showed the effectiveness of adaptively applying Selective Routing. In addition to measuring the performance with the original ordering, we report the average performance for all possible (4!-1) re-orderings.
4.3 Further comparison with PINE
We additionally conduct experiments on a subset of benchmarks (LitM and MMLU) on the runtime latency, perplexity, and collision rate of PINE and RoToR, to validate our claims on Sec. 3.2.
@@ 5 Results & Analysis
Overview
We show results for KGQA in Tab. 2, and results for MMLU in Tab. 3. Results for LitM are in Tab. 1, with a visualization in Appendix Fig. 5. We use lexical sorting for RoToR unless stated otherwise.
Impact of removing index bias on LitM
On LitM, results in Tab. 1 show that, while invariant LMs show stable performances regardless of the gold index when index bias is removed (as described in Sec. 4.2) (Appendix Fig. 5(b)), fluctuations are observed if index bias is present (Appendix Fig. 5(a)). This implies that index bias serves as an implicit source of additional positional bias, and invariant LMs benefit from removing it.
@@
We observe that shuffling input segments leads to non-trivial performance degradations in the original model, which exhibits a statistically significant performance drop on our experimented dataset (one-tailed t-test, p < 0.05). In contrast, our proposed RoToR model does not show a statistically significant difference in performance before and after shuffling, indicating that it is more robust against such perturbations (Appendix C). On LitM (Tab. 1), we notice PCW and Set-Based Prompting has impractical performance, with PINE degrading heavily (-4.5%) as number of documents ($k$) increases, while RoToR is less affected. On KGQA (Tab. 2), we show RoToR outperform PINE with lower standard deviation across shuffled segments. The effectiveness increases as N increases (30 -> 50). Practical scalability with increasing $k$ is critical, but we find that previous order-invariant LMs struggle handling larger $k$ (on KGQA and LitM). In contrast, RoToR shows better performance with improved efficiency and robustness.
@@
Experiments against comparing RoToR with PINE (Tab. 4) analyze the following: Runtime: Actual inference times (Appendix Sec. A.6) and find that RoToR outperforms PINE substantially, with efficiency gains increasing alongside $n$. For instance, on LitM (30 docs), RoToR achieves a 43% reduction in total runtime. Perplexity: Lower generation perplexity indicates input representations are closer to in-distribution. On the same LitM dataset, RoToR’s reduced perplexity implies its positional ID assignment effectively mitigates out-of-distribution effects. Collision Rate: PINE’s similarity-based ordering often collides: on average, only 17.3 of 30 similarity values are unique, causing 42% of the segments to be indistinguishable and thus breaking invariance. In contrast, RoToR with lexical sorting only collides if the segment texts are identical. On LitM, this yields zero collisions, preserving full invariance.
@@
Selective Routing on MMLU (Tab. 3) is a representative of a task that involves not only order-invariant, but also order-sensitive (e.g., "None of the above"), inputs. Therefore, single use of order-invariant models does not always outperform the original model, limiting applicability of order-invariant models to practical listwise tasks. For example, we observe significant performance drops for Set-based Prompting in MMLU, falling short of half the performance of the original model on initial ordering. Using RoToR with Selective Routing to handle order-sensitive inputs can be a solution to this case, which outperforms the original model in all possible orderings of candidate choices. Selective Routing improves the generalizability and extends the applicability on practical listwise tasks by adaptively handling order-sensitive inputs.
We additionally report the the RoToR + Selective Routing (Optimal) performance on Tab. 3, which improves by a significant 7% over using a single model. It was evaluated using a relaxed accuracy metric based on the union of predictions from the original and the invariant (RoToR-lexical) model. This highlights the potential of Selective Routing for further accuracy gains through optimizing design choices on routing methods, which we plan to explore in future work.
@@ 6 Conclusion
Our work addresses order-invariance in listwise inputs by identifying core issues in distribution mismatch and adaptive handling of mixed inputs. Our proposed RoToR provides a stable zero-shot order-invariant solution that reduces the complexity of positional ID modification, while MoV adaptively routes between invariant and sensitive LMs to handle real-world scenarios. Together, these methods demonstrate improved performance and reliability on KGQA, LitM, and MMLU benchmarks.
@@ Appendix A Appendix
A.1 Exact values of the lost-in-the middle benchmark
@@
gold idx = 0 4 9 avg. 0
Orig. 66.2% 65.7% 65.7% 65.9% 65.2%
PINE 67.9% 67.8% 67.5% 67.7% 65.9%
RoToR-lex. 69.6% 69.5% 69.3% 69.5% 67.6%
Table 5: Reporting the performance with Llama-3.1-70B-Instruct on a subset of lost-in-the middle benchmark (without index bias) performance. We see that the gains are consistant with the models we mainly investigated (Llama-3.1-8B-Instruct)
Tab. 1 in the main paper reports the full number to the lost in the middle benchmark, and Fig. 5 visualizes them in graph format. Table  5 shows results in different model (Llama-3.1-70B-Instruct).
@@ A.3 Details about preprocessing and evaluation of datasets
All inferences were done with a single NVIDIA RTX A6000 48GB GPU. For reproducibility, we fix the seed and use greedy decoding without any sampling for all experiments. For all datasets we tested, we separate the input to 3 parts: prefix, parallel contexts, and suffix and feed them accordingly to the positional-invariant model. For the case of the original model, we simply join the prefix, context, and suffix text to make one sequential text. For testing with the PCW model, we concat each prefix to the parallel context due to the architectural limitations of the PCW model, which doesn’t have the prefix part (which is processed casually before parallel contexts). For example, on testing the PCW model on MMLU, we append the question and each answer options, which generally result in longer input sequences. For PCW, we use the pcw_generate function, and we additionally utilized the RestrictiveTokensLogitsProcessor provided at the official PCW repository for MMLU classification, to have a similar setup with the log_likelihood option used for other models.
@@ A.3.2 Knowledge Graph Question Answering
For evaluation with Mintaka and WebQSP, we follow the same setup as  Baek et al. (2023). Given the gold answer and model generated answer, the EM score counts if both are exactly the same; Accuracy measures if the generated answer includes the gold answer, and F1 score measures the precision and recall among overlapping words. Since we are testing on a non-trained zero-shot version of the model, we enforce the model to output in json format to make it easier to parse. For the row shuffling setup, we fix the seed to 0, 1, 2 on shuffling rows and report the average scores.
@@ A.6 Replication details on the runtime experiment
Apart from the theoretical runtime efficiency, we measured the actual end-to-end runtime in seconds, to better analyze the practical runtime efficiency between PINE and RoToR. The runtime of each experiment was measured on an ASUS ESC8000-E11 server featuring dual 4th Gen Intel Xeon Scalable processors, 64 CPU threads, 1.1 TB of RAM across 32 DIMM slots, and 8 NVIDIA A6000 GPUs with 48 GB of memory each. Except for the experiments on Llama-3.1-70B-Instruct, we only use a single A6000 GPU for all of the experiments.
@@ A.7 Input data examples
To get a better understanding of the input and output format of the lost-in-the middle dataset and structured knowledge grounding dataset, we provide example inputs for each data format.
@@
Figure 8: Example input for the lost in the middle dataset, without indexing by numbers.
WebQSP Prefix:
<|begin_of_text|><|start_header_id|>system<|end_header_id|>
(Ferdinand Magellan, occupation, seafarer)
(Ferdinand Magellan, country of citizenship, Kingdom of Portugal) (Ferdinand Magellan, date of birth, time: +1480-00-00) (Ferdinand Magellan, instance of, human) (Ferdinand Magellan, date of death, time: +1521-04-27) (Ferdinand Magellan, sex or gender, male) (Ferdinand Magellan, killed by, Lapulapu) Suffix:
Question: what country did magellan represent?, Answer: <|eot_id|><|start_header_id|>assistant <|end_header_id|>
Gold Answer(s):
(‘Reino de Portugal’, ‘Portuguese Kingdom’, ‘Kingdom of Portugal’)
Example generated output:
{"Answer": "Portugal"} (Parsed to: Portugal)
Figure 9: Example input for the WebQSP dataset.
@@ Appendix C Statisticial significance before and after shuffling segments
We have calculated the paired one-tailed t-tests performed on both the baseline (“original”) model and our proposed method (RoToR) at Tab. 8 using the results at Tab. 7 on whether the performance differences between initial ordering v.s. shuffled ordering are statistically significant on the original model and ours. Note that the performance on the Lost-in-the Middle (LitM) was not reported since the LitM benchmark didn’t provide initial ordering information. Specifically, we test whether the mean performance difference, (Before Shuffle - After Shuffle), deviates significantly from zero. As shown in Table 8, the results indicate that while the original model does exhibit a statistically significant drop in performance upon shuffling, our RoToR method does not, suggesting increased robustness to segment-order perturbations.  [^8] The calculation were validated using the online t-test calculator. https://www.mathportal.org/calculators/statistics-calculator/t-test-calculator.php
@@
Before Shuff. After Shuff. Diff. Before Shuff. After Shuff. Diff.
MMLU 68.3 65.5 2.8 68.5 65.7 2.8
WebQSP, ndoc 30 57.6 57.0 0.6 58.8 58.7 0.1
WebQSP, ndoc 50 57.5 56.5 1.0 58.8 58.7 0.1
Mintaka, ndoc 30 48.7 47.9 0.8 51.2 50.9 0.3
Mintaka, ndoc 50 48.6 48.1 0.5 50.8 50.7 0.1
@@
Degrees of Freedom 4
Mean Difference 1.14 0.68
t-Statistic 3.01 1.43
Critical Value 2.132
@@
Table 8: Paired one-tailed t-test results comparing the original model and ours.
@@
Let the five paired differences (Before -- After) be $\{d_{1},d_{2},d_{3},d_{4},d_{5}\}$. Mean Difference: $\bar{d}=\tfrac{1}{5}\sum_{i=1}^{5}d_{i}$. In this case, $\bar{d}\approx 1.14\%.$ Sample Standard Deviation: $s_{d}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(d_{i}-\bar{d})^{2}}\approx 0.8476.$ Standard Error (SE): $\mathrm{SE}=\tfrac{s_{d}}{\sqrt{n}}\approx 0.379.$ t-Statistic: $t=\tfrac{\bar{d}}{\mathrm{SE}}\approx 3.01,\ (df=4).$ Since the critical value at $df=4$ and $\alpha=0.05$ is $2.132,$ we have $3.01>2.132$ Therefore, the difference is statistically significant.
@@
Under the same procedure, $\bar{d}\approx 0.68\%,\ s_{d}\approx 1.0628,\ \mathrm{SE}\approx 0.475.$ t-Statistic: $t\approx 1.4307.$ Since $1.4307<2.132,$ there is no significant difference in performance before and after shuffling for RoToR.

### v2

@@ 4.1 Baselines
Original causal LM with no modifications (Orig.) were compared, which processes text sequentially. Also, we compare RoToR against previous zero-shot order-invariant LMs discussed in Sec. 3.1, namely PCW Ratner et al. (2023), PINE Wang et al. (2024), and Set-Based Prompting McIlroy-Young et al. (2024). We use the LLaMA 3.1 AI (2024) 8B-Instruct444meta-llama/Meta-Llama-3.1-8B-Instruct, Qwen1.5-4B-Chat and Qwen1.5-7B-Chat [^5] Qwen/Qwen1.5-4/7B-Chat as our backbone model for all of our experiments. As our method doesn’t need training, a single A6000 GPU was sufficient to run all of the experiments except for the Llama-3.1-70B-Instruct model (Appendix, Tab. 6). We also conduct experiments on a subset of benchmarks (LitM and MMLU) on the runtime latency, perplexity, and collision rate of PINE and RoToR, to further validate our claims on Sec. 3.2.
@@ 4.2 Benchmarks with listwise inputs
We experiment with three benchmarks involving real-world listwise input data. Examples of exact inputs and outputs are provided in Appendix A.7. All reported scores are rounded to the nearest tenth, except for the standard deviation (rounded to the second decimal place).
Knowledge Graph QA (KGQA) In KGQA tasks, the model takes facts over knowledge graphs represented as (subject, relation, object), and answers the given question based on the given facts. We basically follow the KGQA dataset preprocessing and evaluation setup from  Baek et al. (2023), which uses Mintaka Sen et al. (2022) with Wikidata for knowledge source, and use the Exact Match (EM), Accuracy (Acc), and F1 score metric for evaluation. We also use MPNet Song et al. (2020) as a dense retriever to retrieve top-k facts over each question, and experiment with segment size of 30 and 50. Replication details and example dataset format are at Appendix Sec. A.3 and Fig. 10. Along with measuring the performance of the initial input ordering, we report performance after we shuffle the order of segments with 3 different seeds to see shuffle robustness.
Lost in the middle (LitM) We use the Lost in the Middle (LitM) benchmark Liu et al. (2024), which draws from 2655 queries in the Natural Questions (NQ) dataset. It provides sets of (10, 20, 30) passages, placing the gold passage at predetermined positions (e.g., 0, 4, 9) and filling the remaining slots with irrelevant passages. Following  Liu et al. (2024), the best_subspan_em metric is used. Experiments on LitM found that eliminating the effect of index bias is another important detail for measuring true order robustness: (Appendix Sec. B). Thus, we report experiments with index bias eliminated. The exact prompt and full results including index bias is reported at Appendix Fig. 8 and Sec. A.1.
MMLU The Massive Multitask Language Understanding (MMLU) benchmark Hendrycks et al. (2021) (prompts at Appendix Fig. 11) consists of 57 diverse sub-tasks with a total of 14,015 queries to measure general performance of LMs about the knowledge of the world. Despite its popularity, many works report that performance fluctuates heavily depending on the order of choices Gupta et al. (2024); Pezeshkpour and Hruschka (2023); Wei et al. (2024); Alzahrani et al. (2024); Zheng et al. (2024a) and is widely investigated to measure the positional bias of the model. We notice that a lot of proportions consist of ordering-sensitive inputs, which showed the effectiveness of adaptively applying Selective Routing. We additionally report the average performance for all possible (4!-1) re-orderings.
@@ 5 Results & Analysis
We report results for KGQA in Tab. 2, and results for MMLU in Tab. 3. Results for LitM are in Tab. 1, with a visualization in Appendix Fig. 5. We use lexical sorting for RoToR unless stated otherwise.
@@
We observe that shuffling input segments leads to non-trivial performance degradations in the original model, which exhibits a statistically significant performance drop on our experimented dataset (two-tailed t-test, p < 0.05, Appendix C). In contrast, our proposed RoToR model does not show a statistically significant difference in performance before and after shuffling, indicating that it is more robust against such perturbations. On LitM (Tab. 1), we notice PCW and Set-Based Prompting has impractical performance, with PINE degrading heavily as number of documents ($k$) increases, while RoToR is less affected. On KGQA (Tab. 2), we show RoToR outperform PINE with lower standard deviation across shuffled segments, consistent with different model architectures.
@@
Experiments against comparing RoToR with PINE (Tab. 4) analyze the following: Runtime, scalability: Actual inference times (Appendix Sec. A.6) find that RoToR outperforms PINE substantially, with efficiency gains increasing alongside $n$. For instance, on LitM (30 docs), RoToR achieves a 43% reduction in total runtime. Practical scalability with increasing $k$ is critical, but we find that previous order-invariant LMs struggle handling larger $k$ (on KGQA and LitM). In contrast, RoToR shows better performance with improved efficiency and robustness. Perplexity: Lower generation perplexity indicates input representations are closer to in-distribution. On the same LitM dataset, RoToR’s reduced perplexity implies its positional ID assignment effectively mitigates out-of-distribution effects. Collision Rate: PINE’s similarity-based ordering often collides: on average, only 17.3 of 30 similarity values are unique, causing 42% of the segments to be indistinguishable and thus breaking invariance. In contrast, RoToR with lexical sorting only collides if the segment texts are identical. On LitM, this yields zero collisions, preserving full invariance.
@@
MMLU (Tab. 3) is a representative of a task that involves not only order-invariant, but also order-sensitive (e.g., "None of the above"), inputs. Therefore, single use of order-invariant models does not always outperform the original model, limiting applicability of order-invariant models to practical listwise tasks, i.e., we observe significant performance drops for Set-based Prompting in MMLU, falling short of half the performance of the original model on initial ordering. However, using RoToR with Selective Routing to handle order-sensitive inputs outperforms, or is at least competitive as the original model in all possible orderings of candidate choices. Selective Routing improves the generalizability and extends the applicability on practical listwise tasks by adaptively handling order-sensitive inputs. The RoToR + Selective Routing (Optimal) performance on Tab. 3 was evaluated using a relaxed accuracy metric based on the union of predictions from the original and the invariant (RoToR-lexical) model. This improves significantly, which highlights the potential of Selective Routing for further accuracy gains through optimizing design choices on routing methods, which we plan to explore in future work.
@@ 6 Conclusion
Our work addresses order-invariance in listwise inputs by identifying core issues in distribution mismatch and adaptive handling of mixed inputs. Our proposed RoToR provides a stable zero-shot order-invariant solution that reduces the complexity of positional ID modification, while Selective Routing adaptively routes between invariant and sensitive LMs to handle real-world scenarios. Together, these methods demonstrate improved performance and reliability on LitM, KGQA, and MMLU benchmarks.
@@ Appendix A Appendix
A.1 Full results on the Lost in the Middle Benchmark
@@
Total ndoc (segments) 10 20 30
Gold idx at: 0 4 9 avg. 0 4 9 14 19 avg. 0 4 9 14 19 24 29 avg.
Indexing bias present
Indexing bias removed (main paper)
[…]
@@
gold idx = 0 4 9 avg. 0 4 9 14 19 avg.
Orig. 66.2 65.7 65.7 65.9 65.2 64.3 65.0 66.2 64.8 65.2
PINE 67.9 67.8 67.5 67.7 65.9 65.7 65.9 65.8 65.5 66.1
RoToR-reranking 68.9 69.0 68.8 68.9 67.5 67.5 67.7 67.5 67.6 67.8
RoToR-lexical 69.6 69.5 69.3 69.5 67.6 67.8 67.8 67.7 67.9 68.0
Table 6: Reporting the performance with the Llama-3.1-70B-Instruct model on a subset of lost-in-the middle benchmark (without index bias) performance. We see that the gains are consistant with the models we mainly investigated (Llama-3.1-8B-Instruct)
Impact of removing index bias on LitM
Tab.5 presents the full results on the Lost in the Middle (LitM) benchmark, comparing scenarios where indexing bias is present versus removed. Fig.5 provides a visual representation of these results.
As shown in Appendix Tab.5, invariant LMs exhibit stable performance regardless of the gold index, especially when index bias is removed (as described in Sec.4.2; see also Appendix Fig.5(b)). However, when index bias is present, performance fluctuations are observed (Appendix Fig.5(a)). Notably, RoToR achieves the highest performance across all setups, demonstrating its effectiveness in mitigating positional bias in a zero-shot setting while maintaining overall performance.
These findings suggest that index bias acts as an implicit source of additional positional bias and that invariant LMs benefit significantly from its removal.
Results on other models.
Table  6 shows results in different model (Llama-3.1-70B-Instruct). We can see that the trend continues for the 70B model, implying the robustness of RoToR.
@@ A.3 Details about preprocessing and evaluation of datasets
All inferences were done with a single NVIDIA RTX A6000 48GB GPU. Note that all of the baseline models including our method can be applied directly in a zero-shot, training-free manner. For reproducibility, we fix the seed and disabled random sampling (i.e., used greedy decoding), set the maximum number of new generated tokens to 500, and set the pad_token_id to the same value as the eos_token_id for all experiments. For all datasets we tested, we separate the input to 3 parts: prefix, parallel contexts, and suffix and feed them accordingly to the positional-invariant model. For the case of the original model, we simply join the prefix, context, and suffix text to make one sequential text. For testing with the PCW model, we concat each prefix to the parallel context due to the architectural limitations of the PCW model, which doesn’t have the prefix part (which is processed casually before parallel contexts). For example, on testing the PCW model on MMLU, we append the question and each answer options, which generally result in longer input sequences. For PCW, we use the pcw_generate function, and we additionally utilized the RestrictiveTokensLogitsProcessor provided at the official PCW repository for MMLU classification, to have a similar setup with the log_likelihood option used for other models.
@@ A.3.2 Knowledge Graph Question Answering
For evaluation with Mintaka, we follow the same setup as  Baek et al. (2023). Given the gold answer and model generated answer, the EM score counts if both are exactly the same; Accuracy measures if the generated answer includes the gold answer, and F1 score measures the precision and recall among overlapping words. Since we are testing on a non-trained zero-shot version of the model, we enforce the model to output in json format to make it easier to parse. For the row shuffling setup, we fix the seed to 0, 1, 2 on shuffling rows and report the average scores.
@@ A.6 Replication details on the runtime experiment
Apart from the theoretical runtime efficiency, we measured the actual end-to-end runtime in seconds, to better analyze the practical runtime efficiency between PINE and RoToR. The runtime of each experiment was measured on an ASUS ESC8000-E11 server featuring dual 4th Gen Intel Xeon Scalable processors, 64 CPU threads, 1.1 TB of RAM across 32 DIMM slots, and 8 NVIDIA A6000 GPUs with 48 GB of memory each. We Except for the experiments on Llama-3.1-70B-Instruct, we only use a single A6000 GPU for all of the experiments.
@@ A.7 Input data examples
To illustrate the input and output formats used in our experiments, we provide example inputs for the Lost-in-the-Middle (LitM), Knowledge Graph Question Answering (KGQA), and MMLU datasets. For experiments using the Qwen-Chat model, special tokens were adjusted accordingly. While the example prompts are based on the Llama-3.1-8B-Instruct model, the specific differences in token usage for the Qwen-Chat variants can be observed by comparing the prompts in Fig. 8 and Fig. 9. This adjustment is consistently applied across all datasets. Note that no special tokens are added for the MMLU benchmark, which aligns with the lm-evaluation harness setup.
@@
Figure 8: Example input for the lost in the middle dataset, without indexing by numbers. Prompt for the Llama-3.1-8B-Instruct model.
lost in the middle no indexing (Qwen variant) Prefix:
<|im_start|>system
You are a helpful, respectful and honest assistant. Always answer as helpfully as possible, while being safe. Please ensure that your responses are socially unbiased and positive in nature. If a question does not make any sense, or is not factually coherent, explain why instead of answering something not correct. If you don’t know the answer to a question, please don’t share false information.<|im_end|><|im_start|>user Write a high-quality answer for the given question using only the provided search results (some of which might be irrelevant). Parallel texts:
[Document Title: Thorax] when deep breaths are attempted. Different people …
[Document Title: Chest pain] present with chest pain, and carry a significantly higher … Suffix:
Question: for complaints of sudden chest pain patients should take a<|im_end|>
<|im_start|>assistant
Figure 9: Example input for the lost in the middle dataset, without indexing by numbers, prompt for the Qwen1.5-Chat model.
@@
MMLU Prefix:
The following are multiple choice questions (with answers) about moral disputes.
Norcross agrees that if a being is incapable of moral reasoning, at even the most basic level, then it cannot be Parallel texts:
A. a being of value.
B. an object of moral sympathy. C. a moral agent. D. a moral patient. Suffix:
Answer:
Figure 11: Example input for the MMLU benchmark.
@@ Appendix C Statisticial significance before and after shuffling segments
We conducted paired two-tailed t-tests (Table 9) for both the baseline (“original”) model and our proposed method (RoToR), using the results in Table 8. Our goal was to determine whether the performance differences between the initial ordering and shuffled ordering are statistically significant. We excluded the Lost-in-the-Middle (LitM) dataset because it does not provide an initial ordering. Specifically, the tests evaluate whether the mean performance difference (Before Shuffle - After Shuffle) significantly deviates from zero.
For KGQA, we selected the F1 score as the representative metric among the three available, gathering data points from various task configurations and different models. For MMLU, the results are based on our RoToR variant with selective routing. As shown in Table 9, the original model shows a statistically significant drop in performance when the segments are shuffled, while RoToR does not, indicating increased robustness to segment-order perturbations.  [^8] All statistical calculations were validated using an online t-test calculator: https://www.mathportal.org/calculators/statistics-calculator/t-test-calculator.php
@@
Before
Shuff.
After
Shuff.
Diff.
Before
Shuff.
After
Shuff.
Diff.
Mintaka, Llama3.1-8B-Instruct, ndoc=30 51.9 51.0 0.9 54.1 53.8 0.3
Mintaka, Llama3.1-8B-Instruct, ndoc=50 51.7 51.0 0.7 53.6 53.5 0.1
Mintaka, Qwen1.5-4B-Chat, ndoc=30 34.9 34.7 0.2 35.7 35.5 0.2
Mintaka, Qwen1.5-4B-Chat, ndoc=50 35.8 35.0 0.8 36.2 36.1 0.1
Mintaka, Qwen1.5-7B-Chat, ndoc=30 35.4 35.0 0.4 37.7 37.7 0
Mintaka, Qwen1.5-7B-Chat, ndoc=50 35.7 35.5 0.2 38.0 38.0 0
MMLU, Llama3.1-8B-Instruct 68.3 65.5 2.8 68.5 65.7 2.8
MMLU, Qwen1.5-4B-Chat 53.6 52.6 1 53.7 52.6 1.1
MMLU, Qwen1.5-7B-Chat 60.1 58.6 1.5 60.1 58.8 1.3
@@
Degrees of Freedom 8
Mean Difference 0.94 0.66
t-Statistic 3.71 2.23
Critical Value 2.306
@@
Table 9: Paired two-tailed t-test results comparing the original model and ours.
@@
Let the nine paired differences (Before -- After) be $\{d_{1},d_{2},d_{3},...d_{8},d_{9}\}$. Mean Difference: $\bar{d}=\tfrac{1}{9}\sum_{i=1}^{9}d_{i}$. In this case, $\bar{d}\approx 0.9444\%.$ Sample Standard Deviation: $s_{d}=\sqrt{\frac{1}{n-1}\sum_{i=1}^{n}(d_{i}-\bar{d})^{2}}\approx 0.7632.$ Standard Error (SE): $\mathrm{SE}=\tfrac{s_{d}}{\sqrt{n}}\approx 0.2544.$ t-Statistic: $t=\tfrac{\bar{d}}{\mathrm{SE}}\approx 3.7124,\ (df=8).$ Since the critical value at $df=8$ and $\alpha=0.05$ is $2.306,$ we have $3.71>2.306$ Therefore, the difference is statistically significant.
@@
Under the same procedure, $\bar{d}\approx 0.6556\%,\ s_{d}\approx 0.8833,\ \mathrm{SE}\approx 0.2944.$ t-Statistic: $t\approx 2.2265.$ Since $2.2265<2.306,$ there is no significant difference in performance before and after shuffling for RoToR.